the solubility of AgI in water is  2X10^-5MOL/L AT 298 k.then the solubilty of AgI ina .04M CaI2 SOLUTION AT 298 k is

Grade:11

1 Answers

Vikas TU
7 years ago
           AgI     -------->            Ag+             +                 I-
Ksp = s^2 = 2 x 10^-5
 
Ksp for AgI in CaI2,
Ksp’ = s*(s+0.08)
        = s^2 + 0.08s
        = 2 x 10^-5 + root(20) * 10^-3
       = 0.02 x 10^-3 + 4.21*10^-3
       = 4.23 x 10^-3 approx.